### v1

# KOROL: Learning Visualizable Object Feature with Koopman Operator Rollout for Manipulation

Abstract
Learning dexterous manipulation skills presents significant challenges due to complex nonlinear dynamics that underlie the interactions between objects and multi-fingered hands. Koopman operators have emerged as a robust method for modeling such nonlinear dynamics within a linear framework. However, current methods rely on runtime access to ground-truth (GT) object states, making them unsuitable for vision-based practical applications. Unlike image-to-action policies that implicitly learn visual features for control, we use a dynamics model, specifically the Koopman operator, to learn visually interpretable object features critical for robotic manipulation within a scene. We construct a Koopman operator using object features predicted by a feature extractor and utilize it to auto-regressively advance system states. We train the feature extractor to embed scene information into object features, thereby enabling the accurate propagation of robot trajectories. We evaluate our approach on simulated and real-world robot tasks, with results showing that it outperformed the model-based imitation learning NDP by 1.08$\times$ and the image-to-action Diffusion Policy by 1.16$\times$. The results suggest that our method maintains task success rates with learned features and extends applicability to real-world manipulation without GT object states.
Keywords: Manipulation, Koopman Operator, Visual Representation Learning
1 Introduction
Figure 1: left: Vanilla Koopman operators rely on ground-truth state which may be difficult to obtain in real-world settings. right: In contrast, we propose
Humans possess an extraordinary ability to manipulate objects, discerning position, shape, and other properties with just a glance. How can robots be endowed with similar perceptual and dexterous manipulation capabilities? Traditional control and optimization approaches typically require detailed models of the system dynamics [hogan2020tactile, mordatch2012contact]. However, these models can be difficult to derive and often lack the flexibility and generalizability needed to adapt to task or environment changes. End-to-end data-driven methods overcome these challenges by learning actions directly from observations [levine2016end, zhao2023learning, shridhar2023perceiver]. While these methods can make minimal assumptions, they often require a large number of demonstrations to master basic skills due to the high dimensionality of the inputs.
To combine sample efficiency of traditional model-based approaches with high generalizability of deep learning methods, one branch of recent work has focused on learning dynamics models to plan trajectories. These methods embed learning into various models, such as Koopman operator [han2023utility, lusch2018deep], Dynamic Movement Primitives (DMP) [bahl2020neural], Neural Geometric Fabrics [xie2023neural, van2024geometric], and more [nagabandi2018neural], showcasing good performance in simulations. However, they often falter in real-world applications due to their reliance on hard-to-obtain ground-truth (GT) or estimated state information like object poses and contact points. Moreover, the learned dynamics models do not transfer across different tasks without a universal state space design.
We propose an approach to remove the dependency on GT states in model-based manipulation learning (Figure LABEL:fig:teaser). Central to this approach, Koopman Operator Rollout for Object Feature Learning (
2 Related Work
Imitation Learning and Visual Representations for Manipulation.
Imitation learning serves as a primary method for teaching robots to manipulate objects by mapping observations or world states directly to actions. Common approaches include Behavioral Cloning (BC) [pomerleau1988alvinn], Implicit Behavioral Cloning (IBC) [florence2022implicit], Long Short-Term Memory (LSTM) networks [levine2016end, florence2019self], Transformers [zhao2023learning, shridhar2023perceiver], and Diffusion Models [chi2023diffusion]. A significant challenge in imitation learning is representing the visual information of a scene. Strategies include using pre-trained 2D [liu2022instruction] or 3D backbones [shridhar2023perceiver] to output visual embeddings. Other works propose end-to-end learning approaches that simultaneously train the visual encoder and the learning policy [zhao2023learning, chi2023diffusion]. Other techniques focus on learning visual representations through correspondence models [florence2019self], self-supervised novel view reconstruction using Neural Radiance Fields (NeRF) [li20223d, ze2023gnfactor], and Gaussian Splatting [lu2024manigaussian].
Model-Based Learning and Planning.
In robotics, traditional model-based approaches rely on expert knowledge of physics to design system models [mason1985robot, collins2005efficient, mordatch2012contact]. Since traditional methods can miss complex nonlinearities, and end-to-end learning approaches can be data-intensive, a middle ground of data-driven model learning shows promise as a data-efficient way to derive complex models [han2023utility, deisenroth2013gaussian]. Model learning includes a variety of dynamics models such as Koopman operators [korda2018linear, bevanda2021koopman], Deep Neural Koopman operators [lusch2018deep], Dynamic Movement Primitives [bahl2020neural], Neural Geometric Fabrics [xie2023neural, van2024geometric], and others [nagabandi2018neural, nguyen2009model]. Additionally, some studies focus on learning environmental responses to actions to plan a future trajectory [ke2018modeling, yang2023foundation, plate], integrate planning in a generative modeling process [janner2022planning, chi2023diffusion], and seamlessly blend the learning of models with planning [du2019model, chen2023planning].
Koopman Operator Theory.
In the early 1930s, Koopman and Von Neumann introduced the Koopman operator theory to transform complex, nonlinear dynamics systems into linear ones in an infinite-dimensional vector space, using observables as lifted states [koopman1932dynamical, koopman1931hamiltonian]. This transform allows the application of linear system tools for effective prediction, estimation, and control with hand-designed observables [han2023utility, brunton2016koopman, chang2016compositional, bruder2019nonlinear]. Recent methods using neural networks to learn observables have proven more expressive and effective, particularly in chaotic time-series prediction [lusch2018deep, brunton2021modern, yeung2019learning]. Furthermore, the integration of neural network-derived Koopman observables with Model Predictive Control has shown promise in enhancing control tasks [li2019learning, wang2022koopman]. As a significant benchmark, Han et al. [han2023utility] demonstrate the effectiveness of Koopman operators in manipulation tasks using GT object states. Building on this foundation, we extend the application of the Koopman operator to vision-based manipulation tasks in real-world settings by learning object features directly from images
3 Background: Koopman Operator Theory
In this section, we provide a brief background on the Koopman Operator Theory. Consider the evolution of nonlinear dynamics system $\mathrm{x}(t+1)=F(\mathrm{x}(t))$. Given the original state space $\mathcal{X}$, the Koopman Operator $\mathcal{K}$ introduces a lifted space of observables $\mathcal{O}$ using lifting function $g:\mathcal{X}\rightarrow\mathcal{O}$, to transform the nonlinear dynamics system into a linear system in infinite-dimensional observables space as $g(\mathrm{x}(t+1))=\mathcal{K}g(\mathrm{x}(t))$.
In practice, we approximate the Koopman operator by restricting observables to be a finite-dimensional vector space. Let $\phi(\mathrm{x}(t))\in\mathbb{R}^{p}$ represent a finite dimensional approximation of observables $g(\mathrm{x}(t))$, and a matrix $\mathbf{K}\in\mathbb{R}^{p\times p}$ approximate the Koopman operator $\mathcal{K}$. Thus, we rewrite the relationship as
$\phi(\mathrm{x}(t+1))=\mathbf{K}\phi(\mathrm{x}(t)).$ (1)
Given a dataset $D$, in which each trajectory $\tau=[\mathrm{x}(1),\mathrm{x}(2),\cdots,\mathrm{x}(T)]$ containing $T$ time steps, we can learn $\mathbf{K}$ by minimizing the state prediction error [brunton2021modern]
$\mathbf{J}(\mathbf{K})=\sum_{\mathrm{x}\in D}\sum_{t=0}^{t=T-1}\|\phi(\mathrm{x}(t+1))-\mathbf{K}\phi(\mathrm{x}(t))\|^{2}.$ (2)
In manipulation tasks, we define the state $\mathrm{x}(t)=[{\mathrm{x}_{r}(t)}^{\top},{\mathrm{x}_{o}(t)}^{\top}]^{\top}$ to include the robot state $\mathrm{x}_{r}(t)$ and object state $\mathrm{x}_{o}(t)$, as we care about how objects move as a result of robot’s motion. Moreover, since our goal is to minimize the imitation error of the robot state $\mathrm{x}_{r}(t)$, we design observables $\phi(\mathrm{x}(t))$ that include lifted robot and object states as
$\phi(\mathrm{x}(t))=[{\mathrm{x}_{r}(t)}^{\top},\psi_{r}({\mathrm{x}_{r}(t)}),{\mathrm{x}_{o}(t)}^{\top},\psi_{o}({\mathrm{x}_{o}(t)})]^{\top}\;\forall t,$ (3)
where $\psi_{r}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\prime}$ and $\psi_{o}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m\prime}$ are vector-valued lifting functions that transform the robot and object state respectively. We can thus retrieve the desired robot state by selecting the corresponding elements in $\phi(\mathrm{x}(t))$. Let $\phi^{\raisebox{0.60275pt}{$\scriptscriptstyle-1$}}$ denote the unlifting function to reconstruct the robot state from observables, $\mathrm{x}_{r}(t)=\phi^{\raisebox{0.60275pt}{$\scriptscriptstyle-1$}}\circ\phi(x(t))$ (we can also reconstruct the object state $\mathrm{x}_{o}(t)$ in the same way). Considering the lifting function Equation 3, the unlifting function can be represented as
${\mathrm{x}_{r}(t)=\phi^{\raisebox{0.60275pt}{$\scriptscriptstyle-1$}}\circ\phi(\mathrm{x}(t))=[\mathrm{I}_{n\times n},\mathrm{0}}_{n\times(n\prime+m+m\prime)}]\cdot\phi(\mathrm{x}(t)),$ (4)
where $\mathrm{I}_{n\times n}$ and $\mathrm{0}_{n\times n}$ denote an identity matrix and zero matrix respectively. To streamline notation throughout this paper, we define $\hat{\mathrm{x}}_{r}(t+1)=\mathbf{K}^{\prime}(\mathrm{x}_{r}(t),\mathrm{x}_{o}(t))$, where $\mathbf{K}^{\prime}\coloneqq\phi^{-1}\circ\mathbf{K}\circ\phi$.
Figure 2: Training and Execution Pipeline. During training,
4 Method
We propose
5 Experiments
In this section, we evaluate the performance of